Grade 12Physical Chemistry

pka for acetic acid is 4.74.what should be the ratio of concentrations of acetic acid and acetate ions to have a solution with pH 5.74?

lokesh palingi
12 Years agoGrade 12

2 Answers

Sunil Kumar FP
12 Years ago
acidic buffer solution
ph=pka + log[salt/acid]
5.74=4.74 +log[salt/ acid]
ratio of their conc=10

rhanks and regards
sunil kr
askIItianfaculty
Aarti Gupta
12 Years ago
pH = pka + log[acetate]/[acetic acid] 5.74 = 4.74 + log[acetate]/[acetic acid] so, log[acetate]/[acetic acid] = 5.74-4.74 = 1.00 thus, [acetate]/[acetic acid] = 0.011 Thanks & regards Aarti Gupta askiitians Faculty